OpenStudy (anonymous):

Rewrite the slope intercept equation of the line y = 1/3x - 2 in standard form.

OpenStudy (anonymous):

Standard point slope form is \[y - y^1 = m (x - x^1)\]  \[y + 2 = 1/3x - 0\]